OpenStudy (anonymous):

write an equation for a line that passes through (4, -12) and is parallel to 4x+3y=12.

OpenStudy (katrinakaif):

For parallel equations, we have to use the same slope as the given equation. In your equation, 4x + 3y = 12 convert it into y = mx + b format: 3y = -4x + 12 y = -4/3x + 4 We know the slope is -4/3 Now use teh equation: (y-y1) = m (x-x1) But substituting the given values: (x1,y1) = (4,-12) and m = -4/3

OpenStudy (anonymous):

thank you (; again,.. :p

OpenStudy (katrinakaif):

Haha anytime ;)